OpenStudy (melissa_something):

Substituting back into the original equation question:

OpenStudy (melissa_something):

This is the question, I got x= 2, 12

OpenStudy (melissa_something):

My question is how could I make checking them easier on the test :/

OpenStudy (anonymous):

im actually not sure sorry!!

OpenStudy (melissa_something):

It's okay @Ethanwestbrooks

zepdrix (zepdrix):

Make checking them easier? To determine if any solutions are extraneous?\[\large\rm \log_2(-number)=bad\]Easy as that. You don't need to do any further calculations. x=2 is extraneous.

OpenStudy (melissa_something):

I thought it was only negative numbers and 0 that can't be true? How do you know?

zepdrix (zepdrix):

Correct, the range of your log function is only positives, no zero, no negative.\[\large\rm \color{red}{\log_2(2-6)}+\color{red}{\log_2(2-4)}-\log_2(2)\]We're getting log of negative value in both of these red spots, ya?

zepdrix (zepdrix):

Sorry sorry, I mean the domain of your log function, the values that can be `put inside of it`.

zepdrix (zepdrix):

I colored them in red, do you not see the color? The first one is log(-4) while the other is log(-2). -4 and -2 are outside of the values we're allowed to put into the log function.

OpenStudy (melissa_something):

If you sent a picture or drawing I didn't get anything :(

zepdrix (zepdrix):

When I do this: \(\large\rm \color{red}{\log_2(2-6)}+\color{red}{\log_2(2-4)}-\log_2(2)\) is the fancy math not showing up? :) You must be using a cellphone or some nonsense XD lol

OpenStudy (melissa_something):

I refreshed and it worked! Lol

OpenStudy (melissa_something):

I see where you got the log (-4) but wouldnt the other one be log (-6) ?

zepdrix (zepdrix):

2-6 = -6+2 = -4 this value becomes slightly `less negative` when we add positive 2. 2-4 = -4+2 = -2 again, becomes slightly less negative.

OpenStudy (melissa_something):

Oh my gosh, thank you. I get it. So did I just get lucky with this problem or do other problems require you to do it again?

zepdrix (zepdrix):

Do what again? :o Plug the other values in?

OpenStudy (melissa_something):

Yes, like for x=12, would I have to go back and do the same?

zepdrix (zepdrix):

Yes you should plug in each solution and see if they hold true to the original equation:\[\large\rm \log_2(12-6)+\log_2(12-4)-\log_2(12)\]This gives us:\[\large\rm \log_2(+)~+~\log_2(+)-\log_2(+)\] We don't get any negative funny business from x=12, so we can finally conclude,\[\large\rm x=12\]\[\large\rm x\ne2\]

OpenStudy (melissa_something):

I got it, thank you so much @zepdrix <3

zepdrix (zepdrix):

yay c:

OpenStudy (melissa_something):

You are a life saver! :D